How many electrons will have the charge of one coulomb?

Grade:12

4 Answers

anuj Yadav
13 years ago

 

Charge on one electron: 1.60 × 10-19C

 Number of electrons to have charge of 1C =  

SAGAR SINGH - IIT DELHI
13 years ago

Dear krishna,

Two bodies with the same type of electricity repelled each other, and two bodies charged with di erent types of electricity attracted each other. We state this as an electrical law, that like charges repel each other and unlike charges attract each other.
The charge carried by the glass rod that has been rubbed with silk is called a positive charge, and the charge on the rubber rod that has been rubbed with wool or fur is called a negative charge.

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 If you like this answer please approve it....

win exciting gifts by answering the questions on Discussion Forum

Sagar Singh

B.Tech IIT Delhi

Rohit Singh
13 years ago

1 electron=1.6×10-19   coulomb

1 coulomb=1/1.6×10-19    Electrons  =6.25×1018 electrons     


shivaraju r
13 years ago

6.25*10^18